A certain college has 52 students participating in games.30 play netball, 25 play football , 10 play handball and football ,13 play netball and handball,9 play netball and football while 2 play all the three games. How many students play handball and no other game?  




1
Expert's answer
2021-02-24T07:50:12-0500

Draw Venn diagram


"n(Nb\cup Fb\cup Hb)=52"

"n(Nb)=30, n(Fb)=25"

"n(Hb\cap Fb)=10, n(Nb\cap Hb)=13,"

"n(Nb\cap Fb)=9, n(Nb\cap Hb\cap Fb)=2"

"n(Nb\cup Hb\cup Fb)=n(Nb)+n(Hb)+n(Fb)"

"-n(Nb\cap Hb)-n(Nb\cap Fb)-n(Hb\cap Fb)"

"+n(Nb\cap Hb\cap Fb)"

"52=30+n(Hb)+25-10-13-9+2"

"n(Hb)=27"

"27-11-2-8=6"

6 students play handball and no other game.